OpenStudy (anonymous):

Help?!!! Solve 4 csc x + 6 = -2 on the interval : [0 ,2pi)? A. pi/3, 5pi/3 B. pi/6, 5pi/6 C. 2pi/3, 4pi/3 D. 7pi/6, 11pi/6

OpenStudy (anonymous):

simplify the equation first\[4 \csc x + 6 = -2\]\[4 \csc x = -8\]\[\csc x = -2\]\[\frac{1}{\sin x} = -2\]\[\sin x = -\frac{1}{2}\]can you continue ?

OpenStudy (anonymous):